Salami and more deli sells a 6-foot sandwich for parties. It weighs 8 pounds. Assume the weight per foot is constant.
a. How much does a sandwich 0-feet long weigh ? (And I mean zero)
b.write a proportion to find the length of a 12 -pound sandwich.
(Please help I would love u forever lol )

Answer by stanbon(75887)

Salami and more deli sells a 6-foot sandwich for parties. It weighs 8 pounds. Assume the weight per foot is constant.
a. How much does a sandwich 0-feet long weigh ? (And I mean zero)
Ans: 0 lbs
-------------------
b.write a proportion to find the length of a 12 -pound sandwich.
length/lbs = length/lbs
----
L/12 = 6/8
----
Length = 12(6/8)
Length = 9 ft
===============
Cheers,
Stan H.
------------------